A husband buys a helium-filled anniversary balloon for his wife. The balloon has a volume of 3.0L in the warm store at 75?F. When he takes it outside, where the temperature is 46?F, he finds it has shrunk.

By how much has the volume decreased? (answer in ?V= L)
Answer #1

using a charles law,

the pressure is constant, the volume is directly proportional to the temperature.

V1 / T1 = V2 / T2                                    { eq. 1 }

where, T1 = 750 F     & T2 = 460 F

temperature change into Rankin

T1 = (75 + 460 ) = 535

T2 = (46 + 460) = 506

V1 = 3 L            and          V2 = ?

insrting all these values in eq. 1,

V2 = V1 x T2 / T1

V2 = (3 L) (506) / (535) = 2.837 L

the volume has decreased by,

(V1 - V2) = (3 - 2.837)L = 0.163 L